prettyariannaher11:

can you show me how i got this x at most 8

anthony5012005:

whats the problem

prettyariannaher11:

i wanna how im suppose solve it

anthony5012005:

well I'm going to try but what is the question we r dealing with

prettyariannaher11:

this a long one

Sailor:

u can screenshot it so u can save time from typing it

prettyariannaher11:

the area of a rectangle dog run can be no more than 120 square feet if the length of the dog run is 15 feet which inequality represents all the possible lengths of the bases

prettyariannaher11:

i dont know how to screen shot

Sailor:

o ok

anthony5012005:

Length x width= answer

anthony5012005:

so 15 x 8 =

surjithayer:

\[15\times x \le 120\] \[15x \le 120\] \[x \le \frac{ 120 }{ 15 }\] \[x \le 8\]